OpenStudy (anonymous):

last problem, i swear *x+6*-*x-5*=1 The stars mean its in a radical

OpenStudy (anonymous):

\[\sqrt{x+6}-\sqrt{x-5}=1\]

OpenStudy (anonymous):

yes thats it

OpenStudy (anonymous):

The answer is x = 30 (I knew 30+6 is 36 and the square root of that is 6 and 30-5 is 25 and the square root of that is 5 so 6-5 = 1)

OpenStudy (anonymous):

how did you get that

OpenStudy (anonymous):

I used guess and check on this one :P. but give me a moment and I may be able to figure out how to go through the steps to prove it.

OpenStudy (anonymous):

Sigh. This one takes a little while to type out in LaTex :P Hold on..

OpenStudy (anonymous):

sqrt ( x+6) = 1 + sqrt (x+5)

OpenStudy (anonymous):

square both sides

OpenStudy (anonymous):

x+6 = 1 + 2 sqrt (x+5) + (x+5)

OpenStudy (anonymous):

x+6 - 1 - (x+5) = 2 sqrt (x+5)

OpenStudy (anonymous):

err

OpenStudy (anonymous):

sqrt ( x+6) = 1 + sqrt (x- 5)

OpenStudy (anonymous):

x+6 = 1 + 2 sqrt ( x - 5) + x- 5 , when you square both sides

OpenStudy (anonymous):

6 + 5-1 = 2 sqrt ( x - 5)

OpenStudy (anonymous):

10 / 2 = sqrt (x-5)

OpenStudy (anonymous):

25 = x - 5 , x = 30

OpenStudy (anonymous):

THIS IS A JOKE

OpenStudy (anonymous):

anwar, we are trying to do the proof

OpenStudy (anonymous):

in my post

OpenStudy (anonymous):

thanks for the help cantor, I was looking at conjugates for some reason

OpenStudy (anonymous):

\[\sqrt{x+6} - \sqrt{x-5} = 1\] \[\sqrt{x-5} = \sqrt{x-6} - 1\] squaring both sides \[x - 5 = x - 2\sqrt{x+6} - 7\]\[2\sqrt{x+6} = 12\]\[\implies x = 30\]

OpenStudy (anonymous):

Sorry, third line should have a +7.

OpenStudy (anonymous):

how are you a champion?